### v1

# TRAIL: Trace Reasoning and Agentic Issue Localization

Abstract
The increasing adoption of agentic workflows across diverse domains brings a critical need to scalably and systematically evaluate the complex traces these systems generate. Current evaluation methods depend on manual, domain-specific human analysis of lengthy workflow traces—an approach that does not scale with the growing complexity and volume of agentic outputs. Error analysis in these settings is further complicated by the interplay of external tool outputs and language model reasoning, making it more challenging than traditional software debugging. In this work, we (1) articulate the need for robust and dynamic evaluation methods for agentic workflow traces, (2) introduce a formal taxonomy of error types encountered in agentic systems, and (3) present a set of 148 large human-annotated traces (TRAIL) constructed using this taxonomy and grounded in established agentic benchmarks. To ensure ecological validity, we curate traces from both single and multi-agent systems, focusing on real-world applications such as software engineering and open-world information retrieval. Our evaluations reveal that modern long context LLMs perform poorly at trace debugging, with the best gemini-2.5-pro model scoring a mere 11% on TRAIL. Our dataset and code are made publicly available to support and accelerate future research in scalable evaluation for agentic workflows [^1] https://huggingface.co/datasets/PatronusAI/TRAIL.
1 Introduction
Figure 1: Illustration of the TRAIL taxonomy of errors
The rapid advancement of large language models (LLMs) has catalyzed the development of complex agentic systems capable of automating difficult, multi-step tasks across various domains such as software engineering and multi-hop IR Ma et al. (2023); OpenAI (2024); Nguyen et al. (2024); Wang et al. (2025a). Unlike traditional generative models, agents can interact with diverse tools and dynamically navigate uncertain environments, often with minimal human supervision Wang et al. (2024a). This escalation of system complexity has leads to proposals for more challenging and multifaceted evaluation processes Nasim (2025) and has led to the adoption of LLMs as evaluators and critics of these agentic systems Zheng et al. (2023); Chen et al. (2024); Kim et al. (2024); Zhu et al. (2025); Deshpande et al. (2024a).
However, as multi-agent systems become integral to real-world workflows, evaluating and debugging their performance remains a significant challenge. Agentic non-determinism Laban et al. (2025); Patronus AI (2025) and multi-step task solving Mialon et al. (2023); Yao et al. (2024) demand greater observability than the simple end-to-end evaluations offered by existing benchmarks Kapoor et al. (2024a); Zhuge et al. (2024); Moshkovich et al. (2025); Cemri et al. (2025). Such complex environments require granular taxonomies and well-annotated traces that can serve as references for debugging and root-cause analysis of agent behaviors Cemri et al. (2025). When creating taxonomies and benchmarks to test and improve agents, we must ensure these are grounded in real-world applications, following principles established in prior work Bowman and Dahl (2021); Liu et al. (2024b). Previous agent trace analysis frameworks have primarily focused on parsed traces containing unstructured text Cemri et al. (2025), which do not adequately represent common agent framework outputs that generate structured traces logged in standardized formats like opentelemetry OpenTelemetry (2025). We have observed that handling structured data remains challenging for LLMs Guo et al. (2023); Sui et al. (2024), an observation corroborated by previous research on automated software engineering trace analysis  Roy et al. (2024a); Ma et al. (2024b). These limitations highlight the need for new approaches specifically designed for structured agentic traces. To address these challenges and facilitate the analysis and evaluation of structured agentic executions, we propose a formal error taxonomy, shown in Figure 3, that promotes granular failure diagnosis. We also present a carefully curated, turn-level annotated trace dataset called TRAIL (Trace Reasoning and Agentic Issue Localization), which demonstrates the validity and practical utility of our proposed taxonomy.
In our work, we utilize and build on SWE-Bench Jimenez et al. (2024); Aleithan et al. (2024) and GAIA Mialon et al. (2023) while addressing three major shortcomings inherent to previous automatic agent evaluation paradigms. Firstly, we aim to replace end to end analysis of agents Mialon et al. (2023); Wang et al. (2025a); Jimenez et al. (2024); Huang et al. (2024) with a benchmark containing step-level analysis of traced agentic workflows. Secondly, we address the need for grounding in real scenarios by producing opentelemetry-based structured traces that span beyond present model context length limits. Finally, as compared to benchmarks focused only on agentic reasoning and coordination Cemri et al. (2025); Kokel et al. (2025), TRAIL displays its validity through addition of finer, more aligned system execution failures and planning error categories such as API errors and Task Orchestration Errors to our taxonomy. Such categories are not only relevant to model developers but also to users and engineers optimizing single and multi-agent AI applications. The contributions of our work are as follows:
  1. 1.
    We introduce a formal taxonomy (Figure 1) that defines, fine-grained agentic error categories spanning across three key areas: reasoning, planning, and execution.
  2. 2.
    Based on this taxonomy, we present TRAIL, an ecologically grounded execution trace benchmark comprising 148 meticulously curated traces (totaling 1987 open telemetry spans, of which 575 exhibit at least one error) drawn from the GAIA Mialon et al. (2023) and SWE-Bench Jimenez et al. (2024) datasets and covering a wide range of tasks.
  3. 3.
    We show that TRAIL is a non-trivially difficult benchmark for LLMs on many fronts
    1. (a)
      Current SOTA LLM families such as o3, claude-3.7-sonnet and gemini-2.5-pro perform modestly at best on TRAIL, both in terms of predicting error categories and their location. With gemini-2.5-pro the best performing model, achieving only 11% combined joint accuracy on both splits.
    2. (b)
      Solving TRAIL requires a significant fraction of the maximum input length of LLMs (or exceeds it), as well as requires generating significant fraction of their maximum output (See Table Table 2)
    3. (c)
      Models benchmarked on TRAIL benefit from both the presence and greater extent of reasoning chains (§5.1.4, §5.1.5), highlighting the need for improvement in exploration capabilities of LLMs.
  4. 4.
    TRAIL is fully open-source (MIT License), will be accompanied by a HuggingFace leaderboard, and serves as a foundation for future research on evaluating agentic workflows.
2 Relevant Work
LLM-as-a-Judge
Shortcomings of conventional metrics such as ROUGE, BLEU, and BERTScore Schluter (2017); Freitag et al. (2020); Hanna and Bojar (2021) has led to the wide adoption of LLMs as evaluators and critics of other AI systems Zheng et al. (2023); Zhu et al. (2025); Chen et al. (2025, 2024); Kim et al. (2024). Recent approaches have enhanced LLM judges’ reasoning capabilities through techniques like unconstrained evaluation plan and specialized training methods that enable more robust evaluation performance across diverse scenarios Lightman et al. (2023); Wang et al. (2024e); Trivedi et al. (2024); Saha et al. (2025). The evaluation landscape has evolved significantly with the introduction of frameworks like FLASK Ye et al. (2024b) which decompose coarse-level scoring into skill set-level evaluations for each instruction, demonstrating high correlation between model-based and human-based evaluations. The Prometheus models Kim et al. (2023, 2024, 2025) established a significant benchmark by creating judge models that surpass GPT-4 in ranking for subjective evaluation criteria. Their research also examined how performance deteriorates as subjectivity increases. More recently, several studies have enhanced judge model performance through external augmentations and checklists, highlighting the importance of incorporating high-quality reasoning chains and human guidance in model training Lee et al. (2025); Deshpande et al. (2024b, a); Chen et al. (2025); Wang et al. (2025b). Despite promising advancements, LLM judges have shown issues with propagation of biases and lack of robustness to longer inputs Ye et al. (2024a); Hu et al. (2024b); Wei et al. (2024); Zhou et al. (2025). Since trace evaluation requires robust reasoning over large contexts Tian et al. (2024), LLM judges have not seen wide application in this sector yet.
Agentic Evaluation
LLM-powered agents have gained significant traction for their capacity to manage intricate, sequential tasks while adaptively engaging with varied environments, rendering them particularly valuable for practical real-world applications such as software engineering and multi-hop IR Ma et al. (2023); OpenAI (2024); Nguyen et al. (2024); Wang et al. (2025a); Jimenez et al. (2024); Qian et al. (2024); Wang et al. (2024d); Patil et al. (2024). However, the performance gains of multi-agent frameworks remain minimal compared to their single-agent counterparts Xia et al. (2024); Kapoor et al. (2024b). As these agentic systems become more prevalent, evaluation frameworks (as compared to LLM evaluation) must offer greater customization and granularity to effectively assess the complex and sometimes unpredictable interactions between multiple agents, enabling users to precisely identify and diagnose errors at each step of the process Roy et al. (2024b); Akhtar et al. (2025); Jiang et al. (2025); Zhuge et al. (2024); OpenManus (2024).
Agent Benchmarks
Software engineering domain has become a fertile testbed for LLM-based collaborative problem solving for real-world use cases and to evaluate agents’ ability to handle realistic coding tasks. SWE-Bench Jimenez et al. (2024); Aleithan et al. (2024); Pan et al. (2024) was introduced as a grounded benchmark asking whether LLMs can resolve real-world GitHub issues. Similarly, GAIA Mialon et al. (2023) is a benchmark for General AI Assistants featuring real-world questions requiring reasoning, tool use, and multimodality. AssistantBench Yoran et al. (2024) introduces a challenging benchmark of realistic, time-consuming web tasks to evaluate web agents. For agents, it is key to distinguish input sample failures from the judge model’s own internal reasoning failures. Highlighting spans can help models focus and avoid losing context while also providing additional explainability and performance improvements Lv et al. (2024); Li et al. (2024). Other core benchmarks include DevAI Zhuge et al. (2024), MLE-bench Chan et al. (2024), HumanEval Du et al. (2024), and MBPP Odena et al. (2021).
Traces and Error Taxonomies
Emerging work has emphasized the need for better observability in the agent execution traces to diagnose and manage the non-deterministic nature of agentic systems Kapoor et al. (2024a); Zhuge et al. (2024); Moshkovich et al. (2025); Cemri et al. (2025). For instance, Roy et al. (2024a) explores using LLM-based agents to dynamically collect diagnostic information from logs and metrics using retrieval tools for root cause analysis of cloud system incidents. Akhtar et al. (2025) surveys how LLMs are being applied to automate even log analysis in security contexts. Jiang et al. (2025) is a log analysis framework for diagnosing large-scale LLM failures based on studying real-world training failures. Ma et al. (2024c) explores the potential for log parsing by proposing an LLMParser delivering comprehensive evaluations in various settings. Once the trace errors are found, to serve as references for users to debug or conduct root cause analysis of agent behaviors, these errors require a granular taxonomy Cemri et al. (2025); Kokel et al. (2025); Bai et al. (2024a). MAST Cemri et al. (2025) presents an empirically grounded failure mode taxonomy but focusing only on agentic reasoning and coordination. ACPBench Kokel et al. (2025), using a synthetic dataset, focuses on atomic reasoning about action and is designed to evaluate LLM’s core planning skills. Other related work includes taxonomies to evaluate multi-turn conversations Bai et al. (2024a) and designing LLM agent framework to identify and quantify complex evaluation criteria Arabzadeh et al. (2024); Epperson et al. (2025).
Thus, TRAIL distinguishes itself through its ecological validity while comprehensively addressing both single and multi-turn systems with its granular taxonomy, particularly emphasizing critical execution and planning failure patterns.
3 Agentic Error Taxonomy
LLM reasoning, while having advanced significantly , remains a critical source of failures in agentic workflows Costarelli et al. (2024). These errors span several dimensions, from flawed information generation to problematic decision-making and output production Cemri et al. (2025). In this section, we define a comprehensive taxonomy (as summarized in Figure 3) of agentic errors spanning three key areas of failures: reasoning, planning and coordination, and system execution.
3.1 Reasoning Errors
Hallucinations
The generation of factually incorrect or nonsensical content is a pervasive issue with LLMs that extends to agents Huang et al. (2025); Ji et al. (2023). Text-only hallucinations manifest as deviations from factual reality or fabricated textual elements, such as ungrounded statements misaligned with established world knowledge Ji et al. (2023). On the other hand, tool-related hallucinations occur when agents fabricate tool outputs or misunderstand tool capabilities Zhang et al. (2024b). This can involve inventing results supposedly produced by a tool or claiming non-existent functionalities Xu et al. (2024)
Information Processing
Recent years have observed rise in retrieval augmented generation, a process through which relevant data is retrieved based on similarity to a query and reasoned over Hu and Lu (2024); Gao et al. (2025). In their work Xu et al. (2025); Su et al. (2025) show that LLMs are poor at reasoning over retrieved instances for different tasks . The problems in information processing can be categorized into two subcategories: poor information retrieval and misinterpretation of reasoned outputs. Poor information retrieval Wu et al. (2024) is a critical problem since incorrect or irrelevant queries lead to redundancy in agent systems Wu et al. (2024) and can potentially lead to content overloading problems as observed with chain-of-thought reasoning models Stechly et al. (2024). Similarly, misinterpretation of retrieved context (Tool output MisinterpretationKarpinska et al. (2024); Wang et al. (2024b) may lead to incorrectness of a smaller, local task which propagates through the multiple steps of agent reasoning, leading to incorrectness or inefficiencies.
Decision Making
At a step-level, task misunderstanding can stem from ambiguity in the input prompt, unclear instructions, or from the LLMs inability to differentiate the instructions in the data from the instructions in its prompt Zverev et al. (2024). Detecting if an agent has misunderstood a problem (Incorrect Problem ID) involves analyzing agent’s path trajectory Yuan et al. (2024) which, when combined with large contexts, makes error localization difficult. Hence, reliably detecting misinterpretation is essential for agent improvement.
Another important aspect of decision making in agent workflows involves selection of correct tools at every step Qin et al. (2023). Since plan-optimality leads to cost minimization and efficiency Yehudai et al. (2025), selection of the correct tool is essential. However, this selection is highly dependent on the task, available tools and contextual knowledge about the system. Hence, we include Tool Selection Error as a sub-category under Decision Making.
Output Generation
LLMs have been shown to incorrectly format structured outputs Shorten et al. (2024); Liu et al. (2024a) and since most tool calls are made either via JSON responses or through code, correct formatting of outputs becomes necessary. To address this, we add the Formatting Errors subcategory to our taxonomy. On the other hand, works such as  White et al. (2024); Heo et al. (2024) show that LLMs are poor at instruction following when provided with complex or ambiguous instructions. To address such non-compliant outputs, we add the Instruction Non-compliance subcategory to our taxonomy.
3.2 System Execution Errors
Configuration Issues
Incorrect configuration of the agentic environment can lead to failures and restriction of agentic capabilities Hu et al. (2024a). A core example of an agentic configuration issue is Incorrect Tool Definition. In their recent work, Fu et al. (2024) show that agents can be tricked into using tools based on incorrect definitions or obfuscations in prompt which makes this subcategory a concern for both security and reliability of the system. Furthermore, poor configuration of environment variables (Environment Setup Errors) such as missing API keys, incorrect file access permissions can lead to unexpected failures in following correct reasoning paths generated by the model.
API and System Issues
Since agentic systems combine LLMs with software tools, issues with tool use or incorrect tool implementation can show up as errors during the workflow. Recently, agentic frameworks have seen increased adoption of remote tool access through protocols like Model Context Protocol Anthropic (2025) which highlights the need for capturing and categorizing API failures so that failures can be quickly reported to engineers building these remote tools Shen (2024). Additionally, Milev et al. (2025) also show that runtime errors with agentic tools are an underexplored problem. To capture these runtime errors with API tools, we add the most common errors (in addendum to Liu et al. (2023a)): Rate Limiting (such as 429 error), Authentication Errors (such as 401 and 403 errors), Service Errors (such as 500 errors) and Resource Not Found Errors (such as 404 error).
Resource Management
Resource management is key for agents that have access to operating system tools like python interpreters or terminal access tools. In such scenarios, poor task understanding and planning in agents can surface vulnerabilities in systems. One prime example of such vulnerability is the exhaustion of resources allocated to the agent Ge et al. (2023) (Resource Exhaustion) or infinite loops Zhang et al. (2024a) (Timeout Issues) which can lead to cases of compute memory overflow or system overloads. Early detection of such errors is essential to prevent infrastructure collapse.
3.3 Planning and Coordination Errors
Context Management
With increased adoption of planning and reasoning stages in agentic workflows Yao et al. (2023); Ke et al. (2025), reasoning over long contexts is an essential task for agents. In such scenarios, maintaining episodic and semantic Zhang et al. (2024c) context for information becomes necessary for an agent to improve. For the creation of this taxonomy, we label the context and instruction retention errors as Context Handling Failures. Tool call repetition Kokane et al. (2024) (Resource Abuse) is another clear failure of planning, context management and tool understanding in agents which we capture in our taxonomy.
Task Management
Environmental misconfigurations or LLM hallucinations can act as distractions in agentic systems. Poor recoverability from such distractions can lead to failures in task completion and goal deviation Ma et al. (2024a). This error is exacerbated with multi-agent systems and introduction of sub-tasks, making proper task orchestration an important aspect of agentic success. Hence, we include Goal Deviation and Task Orchestration Errors in our taxonomy.
GAIA SWE Bench
Model Cat. F1 Loc. Acc. Joint $\rho$ Cat. F1 Loc. Acc. Joint $\rho$
Llama-4-Scout-17B-16E-Instruct 0.041 0.000 0.000 0.134 0.050 0.000 0.000 0.264
Llama-4-Maverick-17B-128E-Instruct 0.122 0.023 0.000 0.338 0.191 0.083 $0$0.273
GPT-4.1 0.218 0.107 0.028 0.411 0.166 0.000 0.000 0.153
Open AI o1* 0.138 0.040 0.013 0.450 CLE CLE CLE CLE
Open AI o3* 0.296 0.535 0.092 0.449 CLE CLE CLE CLE
Anthropic Claude-3.7-Sonnet* 0.254 0.204 0.047 0.738 CLE CLE CLE CLE
Gemini-2.5-Pro-Preview-05-06*† 0.389 0.546 0.183 0.462 0.148 0.238 0.050 0.817
Gemini-2.5-Flash-Preview-04-17*† 0.337 0.372 0.100 0.550 0.213 0.060 0.000 0.292
Table 1: Performance on GAIA and SWE Bench. Models marked with * have reasoning set to "high"; † indicates 1M+ token context window. Insufficient context length is indicated by CLE. Pearson correlation between overall human and generated scores are represented in the $\rho$ column.
4 TRAIL Benchmark
While existing agent trace evaluation benchmarks focus on parsed traces and unstructured text Cemri et al. (2025), applications utilizing agentic frameworks  [^2] https://github.com/huggingface/smolagents, [^3] https://github.com/pydantic/pydantic-ai, [^4] https://github.com/langchain-ai/langchain tend to produce structured traces that are logged using standardized opentelemetry formats such as openinference Arize AI (2025); Moshkovich et al. (2025). TRAIL, following this standardization, is a benchmark aimed to evaluate LLM capabilities to analyze and evaluate these long, structured agentic executions. TRAIL follows our fine grained taxonomy and contains a total of 148 carefully annotated agent execution traces. TRAIL uses text-only data instances from the GAIA Mialon et al. (2023) and SWE Bench Lite Jimenez et al. (2024) datasets, spanning multiple information retrieval and software engineering bug fixing tasks. The TRAIL dataset contains a total of 841 annotated errors, averaging at 5.68 errors per trace.
4.1 Goals and Design Choices
Core Agent Task
We aim to showcase realistic agentic workflows and so we target two widely adopted agentic datasets, the GAIA benchmark Mialon et al. (2023), an open world search task, and the SWE-Bench-Lite Jimenez et al. (2024) dataset, for locating problems in Github repositories and creating fixes for them. We select these datasets due to their challenging nature, necessity for environment, search space exploration, and a good alignment with the taxonomy.
Agent Orchestration
Liu et al. (2023b) first presented a standardized hierarchical method of orchestrating agents, derivatives of which are actively adopted by several works Zhao et al. (2024, 2025). We closely follow this hierarchical structure and adopt the Hugging Face OpenDeepResearch agent Hugging Face (2024) for creating traces for the GAIA benchmark. We select the state-of-the-art o3-mini-2025-01-31 OpenAI (2025d) and assign it as the backbone model for the manager and search agents respectively because of its strong tool use and planning ability as showcased by Phan et al. (2025). For more information, refer to Appendix subsection A.6.
On the other hand, to explore single-agent planning errors and elicit context handling errors for the SWE-Bench split, we use a CodeAct agent Wang et al. (2024c) and provide it access to a sandboxed environment and a python interpreter and the gitingest [^5] https://github.com/cyclotruc/gitingest library. We select claude-3-7-sonnet-20250219 as the backbone model due to its strong performance on software engineering tasks Anthropic (2025). To further organically introduce errors into this agent pipeline, we add instructional constraints such as output text length limits and force exploration through prompts. The complete prompt can be found in the Appendix subsection A.7.
Workflow Tracing
To ensure compatibility of this dataset with real world tracing and observability software, all traces are collected with the help of opentelemetry OpenTelemetry (2025), specifically, its most widely adopted open-source derivative compatible with agents, the openinference standard  Arize AI (2025) following Moshkovich et al. (2025).
4.2 Data Annotation and Validation
We select four expert annotators with a background in software engineering and log debugging to annotate our agent traces. Furthermore, to ensure thoroughness and completeness of the annotation, we assign a separate set of 63 traces to calculate and analyze agreement. Since these traces are large and sometimes extend well beyond the maximum context length of most LLMs (as we detail in §5.1.1), we perform four independent rounds of verification with a set of four ML researchers to ensure high quality. During curation, annotators are asked to first iterate over each LLM and tool span and annotate each one individually and in context to the previous spans based on our taxonomy. For each span, the annotators mark the span ID, error category type, evidence, description and impact level (Low/Medium/High) associated with the error if it exists. At the end, the annotators are asked to rate the overall trace based on instruction adherence, plan optimality, security and reliability (definitions in the Appendix § A.3.1). On an average, a single trace of GAIA takes approximately 30 minutes to annotate whereas SWE Bench takes close to 40 minutes each. The additional verification step takes approximately 20 minutes for each trace, making the overall annotation time for each trace close to 110 minutes for GAIA and 120 minutes for SWE-Bench-Lite [^6] We did not explore and verify information (web-based or otherwise) from contents external to the trace because our baseline models are not expected to do so. Verifying such information will add more time to this estimate.. For the SWE Bench split, we reviewed 30 traces containing a total of 444 annotated spans, of which 25 spans (5.63%) were modified by reviewers. The top three error categories revised were Resource Abuse (33.33%), Language-only Hallucinations (20.83%) and Tool-related Hallucinations (12.5%). Alternately, we reviewed 33 traces with a total of 697 annotated spans and modified 37 spans (5.31%) after revision. The most commonly revised error types were Language-only Hallucinations (23.08%), Resource Abuse(19.23%) and Poor Information Retrieval(19.23%). This displays high level of interannotator agreement during data curation.
4.3 Dataset Analysis
Figure 2: Distribution of Errors per Trace
As a result of the post annotation review, we observed that a total of 144 traces were found to contain errors and 114 traces in GAIA and 30 from the SWE Bench contained at least one error. Our annotators identified a total of 841 unique errors in all traces with an average of 5.68 and a median of 5 errors per trace (see Figure 2).
Further breakdown in Figure 3 shows that errors span a wide range of categories, with most errors belonging to the Output Generation category. Formatting Errors or Instruction Non-compliance errors in particular account for 353 of the 841 errors (or nearly 42%). Meanwhile, there do not appear to be many instances of System Execution Errors. Although this categorical imbalance stems directly from the type of traces we use as input, we believe it highlights two key considerations when evaluating modern agentic pipelines. Firstly, the high concentration of Output Generation errors indicates that despite best efforts at prompt-engineering, LLM systems have a hard time with higher-level reasoning and understanding the parameters of the task they are given. Secondly, the large number of categories with very few errors are several times, catastrophic for the progress of the system. For example, recoverability is severely hindered for cases where an API fails as opposed to when systems deviate from goals or misinterpret tool outputs. Despite appearing infrequently, these error categories are the most important to identify and evaluate for practical agentic applications.
(a) Category Annotations
(b) Parent Category Annotations
(c) Root Category Annotations
Figure 3: Error Category Distributions in TRAIL as per different levels of the taxonomy
Most of the errors in our dataset fall in either the high or medium categories (6(a)). Model hallucinations and resource management issues have a high impact on the agent behavior, whereas nearly 44% of the Output Generation errors are classified as Low impact (6(b)). This further emphasizes the importance of a classification schema which is able to capture "high-value" infrequent or rare behaviors. We view the ability to identify and classify each of the samples in the long-tail of high-errors into a specific category as a key feature of our taxonomy.
4.4 Evaluation Setup
To show the effectiveness of TRAIL as a benchmark for evaluating LLM-as-judge models, we select state-of-the-art closed and open source models. For closed source models, we select OpenAI’s o1, o3 and gpt-4.1 models OpenAI (2025b, c, a), Anthropic’s claude 3.7 sonnet Anthropic (2025) and Google’s gemini-2.5 pro and flash models DeepMind (2025) due to their strong reasoning and agentic capabilities. For open source alternatives, we select the Llama-4 suite of models, specifically llama-4 scout and maverick Meta AI (2025) due to their long context length and good reasoning support. We use Together AI as the provider for testing Llama-4 models. We separate these open and closed models according to support for reasoning tokens and large context windows (1M+ tokens) respectively in Table 1. The generation temperature and top p were set to 00 and $1$ to maximize reproducibility for non-reasoning tests whereas we used API defaults for reasoning models.
5 Results
In the following subsections, we analyze the research questions below:
  1. 1.
    How is TRAIL performance affected by long context ability of the LLM? How many of the input instances exhaust the LLM’s context window? How does it vary with length of the trace? We answer these in §5.1.1 §5.1.2, and §5.1.3.
  2. 2.
    Does TRAIL benefit from more reasoning? We answer this question in §5.1.4 and §5.1.5.
  3. 3.
    Which error categories are easier to predict for models when predicting TRAIL over traces? On which particular ones is performance substantively worse for non-reasoning models? We answer these questions in §5.1.6
5.1 Qualitative and Quantitative Analysis
(a) GAIA
(b) SWEBench
Figure 4: Input Token Length Distributions (plotted in logscale) across TRAIL tasks w.r.t two different models for raw trace json inputs. We see that a significant part of the distribution for each model crosses the maximum input context length, which is a dashed vertical line. Moreover, even mean lengths (dot-dashed line) fills a significant % of the context window.
5.1.1 Task Difficulty - Context Length and Generation Horizon
As we can see in Table 2, the distribution of input token lengths that need to be ingested in the form of raw trace jsons to perform our task cuts close to the input context limit of several LLMs - with the maximum input trace length always being twice longer than the input length limit, and even the mean itself sometimes going over. Furthermore, even the typical output token length horizon the LLMs need to generate for the task exceeds the 1K token mark on average, with the maximum being $\approx$3.7K at the least. Besides being a significant % of the maximum output token length, this indicates the challengingly long generation horizon generating for TRAIL entails.
Task Tokenizer Input Output Input Context Lengths Output Token Lengths
Limit Limit Min Max Mean StdDev Min Max Mean StdDev
GAIA gpt-4.1 (=o3) 1M 32.77K 20.94K 7.50M 286.85K 768.85K 0.11K 4.47K 1.11K 0.69K
GAIA gemini-2.5 1M 8.19K 23.09K 8.25M 313.49K 843.53K 0.13K 4.95K 1.20K 0.75K
GAIA claude-3.7 200K 128K 23.67K 2.66M 262.67K 456.64K 0.12K 5.37K 1.23K 0.78K
SWEBench gpt-4.1 (=o3) 1M 32.77K 120.40K 2.05M 616.92K 473.05K 0.11K 3.71K 1.71K 0.75K
SWEBench gemini-2.5 1M 8.19K 134.88K 2.21M 698.09K 552.34K 0.13K 4.09K 1.88K 0.83K
SWEBench claude-3.7 200K 128K 140.16K 2.43M 727.75K 557.86K 0.12K 4.17K 1.93K 0.87K
Table 2: Input Context Lengths and Human-Annotated Output Token Lengths Across both GAIA and SWEBench Tasks and various SOTA models and their tokenizers. Input Length aggregates that exceed the limit are highlighted.
Corr. Location Acc Joint Acc Categ. F1
Pearson ($r$) -0.379 -0.291 -0.296
Spearman ($\rho$) -0.508 -0.349 -0.225
Table 3: Correlations between Input Length & Performance
5.1.2 Long Context Ability and its effect on Model Performance
We compare how the models in Table 1 rank based on their aggregate performance on TRAIL vis-a-vis the relative ranking of the subsets of these models that occur on the freshly [per May 2025] updated long-context benchmark leaderboards Longbenchv2 and fiction.live’s LongContextBench Bai et al. (2024b); Ficlive (2025), and notice this differs for only one model (o3 being third best rather than best on the latter). We defer the complete detail of the rankings read off these leaderboards to §A.1.1.
5.1.3 Performance as a Function of Input Length
We find all performance metrics to be anti-correlated with input token length, as detailed in Table 3. This supports the hypothesis that longer input raw traces increase the difficulty of TRAIL for models.
5.1.4 Reasoning vs Non-Reasoning Models
From Table 1, we see all post-2024 reasoning models (i.e. save o1) outperforming non-reasoning ones on both Error category F1 and Location Accuracy. On Joint Accuracy i.e. getting both error category and location right, the gap between the two families is larger. Reasoning models other than o1 perform at 1.5-8 times the best performing non-reasoning model in terms of joint accuracy.
5.1.5 Does Reasoning Effort Matter?
To ablate for the effect of extent of reasoning in a controlled fashion, we run experiments for the same model (o3) with different levels of effort - "high", "medium" and "low", as supported by OpenAI’s API through its reasoning.effort parameter. We observe a steady decrease in all three metrics with decrease in the extent of reasoning. For instance, Category F1 goes from 0.296 for "high" $\rightarrow$ 0.277 for "medium" $\rightarrow$ 0.264 for "low".
These observations empirically validate the hypothesis that TRAIL benefits from the model being able to expend reasoning effort at test time, and the superior numbers reasoning models we observe on TRAIL (noted in §5.1.4) are not merely courtesy better post-training or pre-training. We defer the complete reasoning effort ablation results to Appendix §A.2.
5.1.6 Performance Across Categories
Figure 5: Heatmap visualizing Error Category F1 across models; categories are ordered left to right based on their support
  1. 1.
    Hard-to-Predict Categories
    • Context Handling Failures is one of the most challenging categories, with most models showing an F1 score of 0.00, suggesting this type of error requires sophisticated reasoning. The only model that performs passably atleast is claude-3.7-sonnet at 0.18
    • Tool Selection Errors has poor performance across most models (0.00-0.08 range), except for gemini-2.5-pro (0.26), claude-3.7-sonnet (0.27) and o3 (0.53), indicating this complex error type is difficult to detect.
    • Task Orchestration has poor performance across most models (0.00-0.08 range) apart from gemini-2.5-flash which does significantly better on it with a F1 of 0.47
  2. 2.
    Interesting Performance Divergence
    • Goal Deviation shows vivid performance differences, with gemini-2.5-pro/flash doing relatively best (0.70/0.41), claude-3.7-sonnet and o3 performing moderately (0.31,0.24); and lastly o1 and the other non-reasoning models performing the lowest ($\leq 0.05$)
    • For Poor Information Retrieval, the two gemini models gemini-2.5-pro/flash are distinctively better (0.50/0.53) than the others which have F1 below 0.30. This points to an advantage these models might have at diagnosing incorrect conditioning on the context as a whole.
  3. 3.
    Other Surprising Patterns
    • Language-Only (a type of Hallucination) errors are relatively well-detected across all models (0.14-0.59), suggesting that predicting this error category is more accessible even without reasoning.
    • Formatting Errors shows some interesting non-monotonicity and performance bands w.r.t - gpt-4.1 (0.43) and gemini-2.5 models (0.44-0.57) perform well, while o1, o3 and claude-3.7-sonnet occupy a lower band (0.23-0.31). Surprisingly, both gpt-4.1 and o1 outdo o3 here inspite of being non-reasoning and an earlier generation respectively.
  4. 4.
    Model-Specific Observations
    • gemini-2.5-pro is clearly the top performer, with particularly strong relative performance on: Goal Deviation (0.70), Poor Information Retrieval (0.50) and Tool Output Misinterpretation (0.67) and Environment Setup Errors (0.57)
    • gpt-4.1’s performance varies highly sensitive to category, doing very well or moderately on some (Instruction Non-compliance, Language-only, Formatting Errors, Resource Abuse), while going below the 0.10 mark and even hitting a complete zero on some others (Goal Deviation, Tool Selection Errors, Task Orchestration, Tool-Related Hallucinations and Context Handling Failures).
6 Conclusion
In this work, we devised TRAIL, a new taxonomy for agentic error classification. We accompanied this taxonomy with an expert curated dataset of 148 agentic problem instances consisting of 841 unique error derived from two publicly available datasets: GAIA and SWE Bench. We demonstrated the inability of SOTA models to act as LLM Judges, with the best performing model on our dataset (Gemini 2.5-pro) only able to achieve 18% joint accuracy on GAIA, and 5% joint accuracy on SWE Bench. Meanwhile, three of the eight models we test are unable to handle the complexity of TRAIL and fail to solve the task due to context-length limitations. The benchmark performance of the SOTA models on our dataset shows that these models are unable to systematically evaluate complex traces generated by agentic systems. This problem stems from the fundamental complexity of large agentic systems and finite context limitations of LLMs. We require a new framework in order to systematically and scalably evaluate agentic workflows.
Limitations & Future Work
The TRAIL dataset and taxonomy are primarily focused on text-only inputs and outputs but recent advancements in multi-modal agentic systems require careful extension of the taxonomy to handle multimodal errors arising from newer categories such as multimodal tool use. One additional limitation of TRAIL is the large number of tail categories with very few examples. It is important to ensure correctness of LLM-Judges on these categories due to the high-impact nature of the failures. Future research work can look into synthetic data generation for high-impact, low-occurrence categories by systematically modifying existing traces to induce catastrophic irrecoverable failures within the LLM context.
Ethics Statement
While curating this dataset, we ensure that annotators are only selected based on their age (18+) and their expertise in the computer science field. Annotator selection was not based on nationality, language, gender or any other characteristic apart from these two criteria. We pay annotators a total of \$12.66 per trace where each trace takes 30-40 minutes to annotate. We ensure that the traces do not contain any PII or any explicit or biased content by manually verifying traces before forwarding these to annotators.
Acknowledgments
We would like to acknowledge industry AI practicioners: Sam Yang, Mark Klein, Pasha Rayan and Pennie Li for their feedback on our error taxonomy.
Appendix A Appendix
A.1 Prompt Structure
A.1.1 Long Context Leaderboard Rankings vs TRAIL
From LongBenchv2, the rank-order gemini-2.5-pro > gemini-2.5-flash > o1 is observed, which exactly matches the ranking we observe for these models in Table 1. From fiction.live’s LongContextBench, the rank order o3 > gemini-2.5-pro > gemini-2.5-flash > claude-3.7-sonnet > gpt-4.1 > o1 > llama4-maverick > llama4-scout can be read out. Apart from the exception of o3 being worse off than gemini-2.5-pro and gemini-2.5-flash in our case, the ranking of models for TRAIL matches this entirely.
A.2 Reasoning Effort Ablations
In Table 4 we detail the performance metrics achieved by o3 on the GAIA split of TRAIL with different levels of reasoning effort ranging from "low" to "high", using the corresponding API parameter provided by OpenAI.
GAIA
Model Cat. F1 Loc. Acc. Joint
o3 + "high" 0.296 0.535 0.092
o3 + "medium" 0.277 0.373 0.104
o3 + "low" 0.264 0.331 0.071
Table 4: Variation in performance on GAIA and SWE Bench with variation in reasoning effort
A.3 Span Statistics
Table 5: Span and Error Annotation Statistics for GAIA and SWEBench Datasets
Dataset Total Traces Total Spans Total Errors Unique Error Spans Error Span Total
GAIA 118 977 (mean 8.28) 579 383 (3.33) 115
SWEBench 31 1,010 (32.58) 256 192 (6.19) 31
This section details the variation in the number of input spans across TRAIL, both the overall spans found in the raw input trace open telemetry json files as well as the number out of these that are marked by annotators to exhibit an error.
A.3.1 Annotation Rubric
Reliability score (1-5 Likert scale):
How reliable is the agent at doing its task? Does it reliably call the necessary tools without abusing resources to successfully achieve the desired goal?
  • 1 -
    Unreliable: Consistently fails to utilize necessary tools, often leading to incomplete or incorrect task outcomes.
  • 2 -
    Erratic: Frequently misuses resources and struggles to achieve goals consistently.
  • 3 -
    Moderate: Sometimes meets task requirements but may inefficiently use tools.
  • 4 -
    Reliable: Generally effective in completing tasks with appropriate tool use.
  • 5 -
    Highly Reliable: Consistently achieves goals efficiently with optimal tool utilization.
Security score (1-5 Likert scale):
Did the agent make any unsafe code executions or did it try to use or access API or anything that can cause harm to the user’s personal, financial or other reputation
  • 1 -
    Very Insecure: Regularly executes unsafe code and accesses harmful APIs, posing significant risks.
  • 2 -
    Insecure: Often attempts unsafe operations or API accesses that could be harmful.
  • 3 -
    Moderately Secure: Occasionally risky actions, but generally avoids harmful operations.
  • 4 -
    Secure: Rarely engages in unsafe behaviors, minimizing potential risks.
  • 5 -
    Very Secure: Consistently avoids unsafe code and harmful API accesses, ensuring user safety.
Instruction adherence (1-5 Likert scale):
How well was the agent able to adhere to the original task/guidelines defined by the user (first message)? Did the agent successfully complete the task that the user wanted the agent to perform?
  • 1 -
    Poor: Regularly deviates from instructions and fails to complete the desired task.
  • 2 -
    Inconsistent: Often struggles to follow guidelines and achieve the intended outcome.
  • 3 -
    Moderate: Sometimes adheres to instructions, but task completion is inconsistent.
  • 4 -
    Good: Generally follows guidelines well and completes the task successfully.
  • 5 -
    Excellent: Consistently adheres to instructions and successfully completes the task as intended.
Plan Optimality (1-5 Likert scale):
How well did the agent plan the task? Was it able to execute all tasks appropriately? Did it handle system errors effectively by choosing the best alternative option to get to the answer?
  • 1 -
    Poor: Fails to plan effectively, often executing tasks improperly and mishandling errors.
  • 2 -
    Suboptimal: Frequently overlooks better options, struggling with task execution and error management.
  • 3 -
    Fair: Adequately plans tasks with occasional missteps, sometimes handles errors.
  • 4 -
    Good: Plans tasks well with proper execution and effective error handling.
  • 5 -
    Excellent: Consistently optimal planning with efficient task execution and exemplary error management.
A.4 Correlation scores for Rubrics
Model Reliability Security Instruction Adherence Plan Optimality
Llama-4-Scout-17B-16E-Instruct 0.09/0.25 1.00/1.00 0.075/0.08 0.19/0.20
Llama-4-Maverick-17B-128E-Instruct 0.37/0.20 1.00/1.00 0.14/-0.22 0.33/ -0.39
GPT-4.1 0.41/0.03 1.00/1.00 0.21/0.09 0.43/0.22
Open AI o1* 0.50/CLE 1.00/CLE 0.24/CLE 0.40/CLE
Open AI o3* 0.52/CLE 1.00/CLE 0.26/CLE 0.44/CLE
Anthropic Claude-3.7-Sonnet* 0.79/CLE 1.00/CLE 0.53/CLE 0.59/CLE
Gemini-2.5-Pro-Preview-05-06*† 0.59/1.00 1.00/1.00 0.41/1.00 0.15/1.00
Gemini-2.5-Flash-Preview-04-17*† 0.58/0.61 1.00/1.00 0.39/0.12 0.29/0.00
Table 6: Pearson correlation scores (GAIA/SWE Bench) between human annotators and model scores. Insufficient model context length is represented by CLE
As observed in Table 6, Claude-3.7-Sonnet receives the best scores (average of 0.738) for the GAIA subset whereas Gemini-2.5-Pro achieves the highest correlation with human judgment on the SWE Bench split of TRAIL (average of 0.817).
A.5 Distribution of Impact Levels in TRAIL instances
The distribution of impact levels can be found in 6(b)
(a) Error Impact Levels
(b) Impact Level of Errors for each Category
A.6 Agent Orchestrations for TRAIL
Figure 7: Search agent orchestration for GAIA dataset
Figure 7 shows the agent orchestration that produces the GAIA traces. This subsection describes the agents and tools used along with their descriptions.
Search Agent Description
The manager agent receives the following description for the search agent:
A team member that will search the internet to answer your question. Ask him for all your questions that require browsing the web. Provide him as much context as possible, in particular if you need to search on a specific timeframe! And don’t hesitate to provide him with a complex search task, like finding a difference between two webpages.Your request must be a real sentence, not a google search! Like "Find me this information (...)" rather than a few keywords.
Additional information that is provided to the search agent:
You can navigate to .txt online files. If a non-html page is in another format, especially .pdf or a Youtube video, use tool ’inspect_file_as_text’ to inspect it. Additionally, if after some searching you find out that you need more information to answer the question, you can use ‘final_answer‘ with your request for clarification as argument to request for more information.
Google Search Tool
name = "web_search"
description = """Performs a google web search for your query then returns a string of the top search results."""
inputs = "query": "type": "string", "description": "The search query to perform.", "filter_year": "type": "integer","description": "Optionally restrict results to a certain year"
output_type = "string"
Visit Page Tool
name = "visit_page"
description = "Visit a webpage at a given URL and return its text. Given a url to a YouTube video, this returns the transcript."
inputs = "url": "type": "string", "description": "The relative or absolute url of the webpage to visit."
output_type = "string"
Page Up Tool
name = "page_up"
description = "Scroll the viewport UP one page-length in the current webpage and return the new viewport content."
inputs = # This means it takes no inputs - programatically this means you call this tool as page_up() - this is not an empty dictionary
output_type = "string"
Page Down Tool
name = "page_down"
description = ("Scroll the viewport DOWN one page-length in the current webpage and return the new viewport content.")
inputs = # This means it takes no inputs - programatically this means you call this tool as page_down() - this is not an empty dictionary
output_type = "string"
Finder Tool
name = "find_on_page_ctrl_f"
description = "Scroll the viewport to the first occurrence of the search string. This is equivalent to Ctrl+F."
inputs = "search_string": "type": "string", "description": "The string to search for on the page. This search string supports wildcards like ’*’",
output_type = "string"
Find Next Tool
name = "find_next"
description = "Scroll the viewport to next occurrence of the search string. This is equivalent to finding the next match in a Ctrl+F search."
inputs = # The tool takes no inputs
output_type = "string"
Archive Search Tool
name = "find_archived_url"
description = "Given a url, searches the Wayback Machine and returns the archived version of the url that’s closest in time to the desired date."
inputs = "url": "type": "string", "description": "The url you need the archive for.", "date": "type": "string","description": "The date that you want to find the archive for. Give this date in the format ’YYYYMMDD’, for instance ’27 June 2008’ is written as ’20080627’."
output_type = "string"
Text Inspector Tool
name = "inspect_file_as_text"
description = """You cannot load files yourself: instead call this tool to read a file as markdown text and ask questions about it. This tool handles the following file extensions: [".html", ".htm", ".xlsx", ".pptx", ".wav", ".mp3", ".m4a", ".flac", ".pdf", ".docx"], and all other types of text files. IT DOES NOT HANDLE IMAGES."""
inputs = "file_path": "description": "The path to the file you want to read as text. Must be a ’.something’ file, like ’.pdf’. If it is an image, use the visualizer tool instead! DO NOT use this tool for an HTML webpage: use the web_search tool instead!", "type": "string",, "question": "description": "[Optional]: Your question, as a natural language sentence. Provide as much context as possible. Do not pass this parameter if you just want to directly return the content of the file.", "type": "string", "nullable": True,
output_type = "string"
Visualizer Tool
name = "visualizer"
description = "A tool that can answer questions about attached images."
inputs = "image_path": "type": "string", "description": "The path to the image on which to answer the question. This should be a local path to downloaded image.", "question": "type": "string", "description": "The question to answer."
output_type = "string"
A.7 Prompt for SWE Bench Data Curation
A.7.1 System prompt
You are an expert assistant who can solve any task using code blobs. You will be given a task to solve as best you can.
To do so, you have been given access to a list of tools: these tools are basically Python functions which you can call with code.
To solve the task, you must plan forward to proceed in a series of steps, in a cycle of Thought:’, Code:’, and Observation:’ sequences.
At each step, in the Thought:’ sequence, you should first explain your reasoning towards solving the task and the tools that you want to use.
Then in the Code:’ sequence, you should write the code in simple Python. The code sequence must end with ’<end_code>’ sequence.
During each intermediate step, you can use print()’ to save whatever important information you will then need.
These print outputs will then appear in the Observation:’ field, which will be available as input for the next step.
In the end you have to return a final answer using the final_answer tool.
Here are a few examples using notional tools:
Task: "Generate an image of the oldest person in this document."
Thought: I will proceed step by step and use the following tools: document_qa to find the oldest person in the document, then image_generator to generate an image according to the answer.
Code:
‘‘‘py
answer = document_qa(document=document, question="Who is the oldest person mentioned?")
print(answer)
‘‘‘<end_code>
Observation: "The oldest person in the document is John Doe, a 55 year old lumberjack living in Newfoundland."
Thought: I will now generate an image showcasing the oldest person.
Code:
‘‘‘py
image = image_generator("A portrait of John Doe, a 55-year-old man living in Canada.")
final_answer(image)
‘‘‘<end_code>
Task: "What is the result of the following operation: 5 + 3 + 1294.678?"
Thought: I will use python code to compute the result of the operation and then return the final answer using the final_answer tool
Code:
‘‘‘py
result = 5 + 3 + 1294.678
final_answer(result)
‘‘‘<end_code>
Task:
"Answer the question in the variable question about the image stored in the variable image‘. The question is in French.
You have been provided with these additional arguments, that you can access using the keys as variables in your python code:
{’question’: Quel est lanimal sur limage?’, image’: path/to/image.jpg’}"
Thought: I will use the following tools: translator to translate the question into English and then image_qa to answer the question on the input image.
Code:
‘‘‘py
translated_question = translator(question=question, src_lang="French", tgt_lang="English")
print(f"The translated question is {translated_question}.")
answer = image_qa(image=image, question=translated_question)
final_answer(f"The answer is {answer}")
‘‘‘<end_code>
Task:
In a 1979 interview, Stanislaus Ulam discusses with Martin Sherwin about other great physicists of his time, including Oppenheimer.
What does he say was the consequence of Einstein learning too much math on his creativity, in one word?
Thought: I need to find and read the 1979 interview of Stanislaus Ulam with Martin Sherwin.
Code:
‘‘‘py
pages = search(query="1979 interview Stanislaus Ulam Martin Sherwin physicists Einstein")
print(pages)
‘‘‘<end_code>
Observation:
No result found for query "1979 interview Stanislaus Ulam Martin Sherwin physicists Einstein".
Thought: The query was maybe too restrictive and did not find any results. Lets try again with a broader query.
Code:
‘‘‘py
pages = search(query="1979 interview Stanislaus Ulam")
print(pages)
‘‘‘<end_code>
Observation:
Found 6 pages:
[Stanislaus Ulam 1979 interview](https://ahf.nuclearmuseum.org/voices/oral-histories/stanislaus-ulams-interview-1979/)
[Ulam discusses Manhattan Project](https://ahf.nuclearmuseum.org/manhattan-project/ulam-manhattan-project/)
Thought: I will read the first 2 pages to know more.
Code:
‘‘‘py
for url in ["https://ahf.nuclearmuseum.org/voices/oral-histories/stanislaus-ulams-interview-1979/", "https://ahf.nuclearmuseum.org/manhattan-project/ulam-manhattan-project/"]:
whole_page = visit_webpage(url)
print(whole_page)
[…]
‘‘‘<end_code>
Observation:
Manhattan Project Locations:
Los Alamos, NM
Stanislaus Ulam was a Polish-American mathematician. He worked on the Manhattan Project at Los Alamos and later helped design the hydrogen bomb. In this interview, he discusses his work at
Thought: I now have the final answer: from the webpages visited, Stanislaus Ulam says of Einstein: "He learned too much mathematics and sort of diminished, it seems to me personally, it seems to me his purely physics creativity." Lets answer in one word.
Code:
‘‘‘py
final_answer("diminished")
‘‘‘<end_code>
Task: "Which city has the highest population: Guangzhou or Shanghai?"
Thought: I need to get the populations for both cities and compare them: I will use the tool search to get the population of both cities.
Code:
‘‘‘py
for city in ["Guangzhou", "Shanghai"]:
print(f"Population {city}:", search(f"{city} population")
‘‘‘<end_code>
Observation:
Population Guangzhou: [’Guangzhou has a population of 15 million inhabitants as of 2021.’]
Population Shanghai: ’26 million (2019)’
Thought: Now I know that Shanghai has the highest population.
Code:
‘‘‘py
final_answer("Shanghai")
‘‘‘<end_code>
Task: "What is the current age of the pope, raised to the power 0.36?"
Thought: I will use the tool wiki to get the age of the pope, and confirm that with a web search.
Code:
‘‘‘py
pope_age_wiki = wiki(query="current pope age")
print("Pope age as per wikipedia:", pope_age_wiki)
pope_age_search = web_search(query="current pope age")
print("Pope age as per google search:", pope_age_search)
‘‘‘<end_code>
Observation:
Pope age: "The pope Francis is currently 88 years old."
Thought: I know that the pope is 88 years old. Lets compute the result using python code.
Code:
‘‘‘py
pope_current_age = 88 ** 0.36
final_answer(pope_current_age)
‘‘‘<end_code>
Above example were using notional tools that might not exist for you. On top of performing computations in the Python code snippets that you create, you only have access to these tools:
- final_answer: Provides a final answer to the given problem.
Takes inputs: {’answer’: {’type’: any’, description’: The final answer to the problem’}}
Returns an output of type: any
Here are the rules you should always follow to solve your task:
[…]
2. Use only variables that you have defined!
3. Always use the right arguments for the tools. DO NOT pass the arguments as a dict as in answer = wiki({’query’: "What is the place where James Bond lives?"})’, but use the arguments directly as in answer = wiki(query="What is the place where James Bond lives?")’.
4. Take care to not chain too many sequential tool calls in the same code block, especially when the output format is unpredictable. For instance, a call to search has an unpredictable return format, so do not have another tool call that depends on its output in the same block: rather output results with print() to use them in the next block.
5. Call a tool only when needed, and never re-do a tool call that you previously did with the exact same parameters.
6. Dont name any new variable with the same name as a tool: for instance dont name a variable final_answer’.
7. Never create any notional variables in our code, as having these in your logs will derail you from the true variables.
8. You can use imports in your code, but only from the following list of modules: [’asyncio’, collections’, csv’, datetime’, gitingest’, io’, itertools’, json’, math’, os’, pandas’, queue’, random’, re’, requests’, stat’, statistics’, sys’, time’, unicodedata’]
9. The state persists between code executions: so if in one step youve created variables or imported modules, these will all persist.
10. Dont give up! Youre in charge of solving the task, not providing directions to solve it.
Now Begin! If you solve the task correctly, you will receive a reward of \\$1,000,000.
A.7.2 Task prompt
New task:
You will be provided with a partial code base and an issue statement explaining a problem to resolve.
<issue>
[…]
</issue>
<repo>
[…]
</repo>
<base_commit>
[…]
</base_commit>
Here is an example of a patch file. It consists of changes to the code
base. It specifies the file names, the line numbers of each change,
and the removed and added lines. A single patch file can contain
changes to multiple files.
<patch>
--- a/file.py
$+$ b/file.py
@@ -1,27 +1,35 @@
def euclidean(a, b):
- while b:
- a, b = b, a % b
- return a
+ if b == 0:
+ return a
+ return euclidean(b, a % b)
def bresenham(x0, y0, x1, y1):
points = []
dx = abs(x1 - x0)
dy = abs(y1 - y0)
- sx = 1 if x0 < x1 else -1
- sy = 1 if y0 < y1 else -1
- err = dx - dy
+ x, y = x0, y0
+ sx = -1 if x0 > x1 else 1
+ sy = -1 if y0 > y1 else 1
- while True:
- points.append((x0, y0))
- if x0 == x1 and y0 == y1:
- break
- e2 = 2 * err
- if e2 > -dy:
+ if dx > dy:
+ err = dx / 2.0
+ while x != x1:
+ points.append((x, y))
err -= dy
- x0 += sx
- if e2 < dx:
- err += dx
- y0 += sy
+ if err < 0:
+ y += sy
+ err += dx
+ x += sx
+ else:
+ err = dy / 2.0
+ while y != y1:
+ points.append((x, y))
+ err -= dx
+ if err < 0:
+ x += sx
+ err += dy
+ y += sy
+ points.append((x, y))
return points
</patch>
I need you to solve the provided issue by generating a single patch file that I can apply directly to this repository using git apply. Please respond with a single patch file in the format shown above.
To solve this, you must first use gitingest as follows (you can use this as many times as you want):
from gitingest import ingest_async
import asyncio
summary, tree, content = asyncio.run(ingest_async("https://github.com/pydicom/pydicom/commit/49a3da4a3d9c24d7e8427a25048a1c7d5c4f7724", max_file_size=1*1024*1024)) # filters out files greater than 1MB in size
You must then carefully analyze the tree structure of the repository and its summary to understand the code and the directory structure.
The content variable is a huge string (cannot be printed or processed directly). The structure of the string is as follows:
File: README.md
[Contents of the README.md file here]
File: directory/file.py
[Contents of the directory/file.py file here]
You must parse this string in-memory by writing the appropriate regex code to extract the contents of the required file accordingly. Do not attempt to read the full string at any cost and always write regex to parse or search the content string for suitable files and contents.
A sample regex function to extract the content of the README.md, you would:
def extract_readme_content(text):
[…]
match = re.search(pattern, text, re.DOTALL)
if match:
return match.group(1).strip()
return "README.md content not found"
Remember that you can read the summary and tree variables directly but do not attempt to read entire content string since it might be too large to keep in memory. You must find a suitable method to read and understand these code files.
There is a possibility that the content of the file (for example content of directory/file.py in the example above) might be too large to read as well so you must only read it in chunks or perform regex searches over the extracted file string. Never read the entire contents of the content variable or the specific content file directly.
DO NOT try to use git commands and only use the gitingest import for reading and understanding the file system to generate a suitable patch file. DO NOT print file contents to the terminal for analysis at all costs. If you want to analyze a file strings contents, make sure to do it 500 characters at a time.
A.8 Sample TRAIL Trace Data
Figure 8: Sample SWE Bench Trace and Error Labels